OpenStudy (anonymous):

I'm having a really hard time figuring out this Algebra 2 problem, I don't understand it at all. Can someone please help? Christi needs your help on the last question of her math homework. The question reads, "How can key features be used to create a sketch of any polynomial function?" Explain to Christi, in complete sentences, what key features are necessary and how they can be used to create the sketch of a polynomial function.

OpenStudy (anonymous):

A polynomial equation is like ax^3 + bx^2 + c = 0 Significant features include how it behaves for x<<0 and x>>0, large negative and positive values of x, and how many zeroes it has, which is how many times it crosses the x axis (y=0) and where the zeroes are. The degree of the polynomial is the largest exponent, here 3, and this tells you that there are up to 3 unique zeroes.

OpenStudy (anonymous):

Thanks, although I already figured it out and just forgot to close the question sorry.

OpenStudy (anonymous):

You are welcome. congratulations.

OpenStudy (anonymous):

Although I do have one more question I really need help with, can you help with that?

OpenStudy (anonymous):

Unfortunately, I should run. If it's a quickie I'll try.

OpenStudy (anonymous):

Ok thanks, I dont really understand what they mean at all in this question. When solving a radical equation, Beth and Kelly came to two different conclusions. Beth found a solution, while Kelly's solution did not work in the equation. Create and justify two situations: one situation where Beth is correct and a separate situation where Kelly is correct.

OpenStudy (anonymous):

The idea is that a "radical equation." which I think means one involving a root (square root, cube root, etc) can have more than one solution, and depending on something, only one or the other will work. They want an example. Well, sqrt(x) = 3 so we have x = 3^2 = 9, but x=-3 would also give (-3)^2=9 although sqrt(-9) gives an imaginary number result, 3i. Must stop. Good luck.

OpenStudy (anonymous):

Awesome thanks! I had no idea what they were talking about until now lol.